OpenStudy (anonymous):

how is factoring used to solve quadratic equations. Need to demonstrate the process with an example to show.

OpenStudy (anonymous):

(x^2)+x-6=0 (x+3)(x-2)=0 x= 3 x =2

OpenStudy (anonymous):

x=-3 x=2*